A force of 10 N gives an object an acceleration of 6 m/s2. What force will give the same object an acceleration of 18 m/s2.
1
Expert's answer
2020-04-27T10:18:21-0400

The Newton's second law states


"ma=F"

Hence,


"F_2=F_1\frac{a_2}{a_1}=(10\rm \:N)\times\frac{18\:\rm m/s^2}{6\:\rm m/s^2}=30\:\rm N."